A motor boat travels 137.5 miles upstream in 2.5 hours. The next day it travels downstream to the starting point in 5.5 hours. What is the speed of the boat, and of the current?
Answer by amarjeeth123(569)

Let the speed of the boat be x
Let the speed of the current be y.
When it travels upstream
x+y=137.5/2.5
x+y=55........equation 1
When it travels downstream
x-y=137.5/5.5
x-y=25........equation 2
Adding 1 and 2 we get,
2x=55+25
2x=80
x=40
y=15
The speed of the boat is 40 miles/hr.
The speed of the current is 15 miles/hr.